# InterActHuman: Multi-Concept Human Animation with Layout-Aligned Audio Conditions

### v1

@@ Appendix A Algorithm of Our Model Implementation
7:for $i\leftarrow 1$ to $N$ do
8:     $T_{i}\leftarrow\text{Rephrase}(X_{i},T)$triangleright Generate detailed prompt via Qwen2.5-VL.
9:     $\mathbf{x}_{i}\leftarrow\text{VAE\_Encoder}(X_{i})$triangleright Encode reference image.
10:     $\mathbf{a}_{i}\leftarrow\text{wav2vec}(Y_{i})$triangleright Extract audio features.
@@
12:$c_{\text{text}}\leftarrow\{T,T_{1},\ldots,T_{N}\}$triangleright Aggregate text conditions.
13:$z_{S}\sim\mathcal{N}(0,I)$triangleright Initialize the noisy video latent.
@@
17:         for $i\leftarrow 1$ to $N$ do
18:              Reference Injection: Inject $\{\mathbf{x}_{i}\}$ via concatnation and self-attention.
@@
20:              $\mathbf{Q}^{v}\leftarrow\text{Proj}_{q}(\mathbf{h}^{v})$.  $[\mathbf{K}^{r}_{i},\,\mathbf{V}^{r}_{i}]\leftarrow\text{Proj}_{k,v}(\mathbf{h}^{r}_{i})$.
@@
31:              $\mathbf{Q}^{v}\leftarrow\text{Proj}_{q}(\mathbf{h}^{v})$.  $[\mathbf{K}^{\text{mute}}_{i},\,\mathbf{V}^{\text{mute}}_{i}]\leftarrow\text{Proj}_{k,v}(\mathbf{a}_{i}^{\text{mute}})$.  $[\mathbf{K}_{i},\,\mathbf{V}_{i}]\leftarrow\text{Proj}_{k,v}(\mathbf{a}_{i})$.
@@
Appendix B Experiment Details of Our Model
B.1 Implementation Details
@@
B.2 Ablation Details
In Tab. 4 and Fig. 5 of the paper, we validate our local audio injection designs via ablation on three variants: 1) Global audio lin2025omnihuman applies audio across the entire feature map; 2) ID Embedding injects audio features with a learnable ID embedding without mask prediction; 3) Fixed Mask uses predefined static spatial masks.
For global audio, it is identical to the pretrained audio-driven model lin2025omnihuman . For fixed mask, it is a manually input rectangle mask upon the audio cross-attention of pretrained audio-driven model lin2025omnihuman . It requires the generated human being static and it needs the user to ensure there is only one person inside a mask.
For ID embedding, here we provide some implementation details. Inspired by detection transformers carion2020end , we introduce a set of $N$ learnable ID embeddings $\mathbf{E}_{query}\in\mathbb{R}^{N\times C}$. These embeddings serve as identity tokens for the $N$ individuals. Specifically, we add the same learnable ID embedding $\mathbf{e}_{query}\in\mathbb{R}^{C}$ to a paired reference image and audio segment. In this way, we expect the model to implicitly match an individual (defined by a reference image) and audio segments in the output video to get synchronized lip movements. For the ID embedding dataset preparation, we incorporate multi-person videos into training. We use Active Speaker Detection (ASD) to extract talking persons’ identities as a bounding box with identity-number. Then we use the bounding box and SAM2 masks to match the identity-number of each timestep of audio based on their IoU. Finally we use paired audio with timestep and reference image to generate video, and use flow-matching loss only to supervised it. No mask information is provided to the model and no mask supervision adopted.
@@ B.2 Ablation Details
Appendix C Details of Audio-driven Base Model’s Architecture and Training
This appendix provides a detailed description of the network structure and training specifics for the pretrained audio-driven single-person video generation model, OmniHuman lin2025omnihuman . As noted in our main paper, the text-to-video base model undergoes post-pretraining for audio conditioning, following the OmniHuman lin2025omnihuman methodology. Subsequently, our multi-concept framework is built and trained on this foundation.
C.1 Audio-driven Base Model
@@
The DiT Blocks are based on the dual-stream Diffusion Transformer (DiT)esser2024scaling . This transformer processes video and text tokens through multiple self-attention layers and feedforward networks (FFNs) to learn representations that are both shared and modality-specific. SwiGLU is employed as the activation function to enhance nonlinear modeling capabilities. Additionally, AdaSinglechen2023pixart is used for efficient timestep modulation. Moreover, two-thirds of the FFN weights in the deeper layers share parameters, creating a hybrid-stream design that preserves model capacity while substantially improving parameter efficiency.
C.2 Mixed Training Strategy
We utilize a multi-condition training strategy based on the framework established in OmniHuman lin2025omnihuman . Our method employs a two-stage progressive training scheme to develop the base model’s capabilities. In the initial stage, the model is trained solely on text-to-video (T2V) data to build fundamental video generation abilities. The second stage introduces audio-synchronized datasets to expand the model’s functionality to include audio-driven generation and reference image injection. For this second stage, we follow two core principles from OmniHuman: 1) Tasks with stronger conditioning can leverage tasks with weaker conditioning and their associated data to broaden the effective training dataset; 2) Tasks with stronger conditioning should be assigned proportionally lower training ratios. Guided by these principles, we first train the reference image injection capability before introducing the audio-driven generation objective. This staged methodology facilitates efficient knowledge transfer while ensuring stable training dynamics throughout the multi-condition learning process.
Appendix D Additional Details for Audio-driven Base Model Dataset
D.1 Dataset Curation
@@
2. Human. Utilizing the annotated video captions, we implement a rule-based system to detect keywords such as "people", "human", "men", "women", "girl", and "boy". If any of these keywords are found, the video is categorized as human-related. This technique ensures the dataset is effectively filtered to comprise only videos pertinent to human activities and interactions.
@@
4. Visual Quality. We utilize Q-align wu2023q to evaluate the visual quality of the videos, filtering out clips that do not meet a predefined threshold. This procedure ensures the dataset maintains a high standard of visual clarity, which is essential for accurate analysis and robust model performance. By eliminating low-quality segments, we enhance the overall dependability and utility of the dataset.
5. Aesthetics. We employ Q-align wu2023q to assess the aesthetic appeal of the videos and discard clips falling below a set threshold. Through aesthetic quality assessment, we can filter out videos containing post-production elements, thereby improving the quality of the training dataset. This measure ensures the dataset is composed of natural and unaltered video content, which better represents real-world scenarios and boosts the robustness and generalization of the trained models.
6. Motion. We use Raft teed2020raft to calculate the optical flow of the videos and filter out clips exhibiting overly intense motion. This step ensures that the dataset includes only video clips with moderate and significant motion, which are more appropriate for analysis and model training. By removing clips with extreme motion, we enhance the stability and quality of the dataset, leading to more dependable results.
7. Syncnet. For audio-driven content, we employ SyncNet chung2017out to determine if the lip movements are synchronized with the audio. Videos displaying considerable asynchronization are excluded. This step ensures that the dataset consists only of high-quality, synchronized audio-visual data. By removing out-of-sync segments, we improve the overall quality and reliability of the dataset. Ultimately, we gather 2,000 hours of audio-driven data.
D.2 Dataset Analysis
@@
  • Human Size. Human size indicates the portion of the human body visible within the video frame. It is categorized into the following levels: Portrait (head and shoulders), Chest (from head to chest), Waist (from head to waist), Knees (from head to knees), and Full Body (the entire body is visible). To determine this, we first detect body keypoints using RTMpose jiang2023rtmpose , and then classify the human size based on the confidence scores of these keypoints. This method ensures a robust and accurate categorization of the visible human body extent in each video.
  • Motion Amplitude. After extracting body keypoints from the video, the amplitude of human motion is computed by measuring the displacement of the chest keypoint over time, relative to the width of the shoulders. Based on these calculations, we classify motion amplitude into four categories: Slight($<0.1$), Moderate($0.1-0.2$), Significant($0.2-0.3$), and Extreme($>0.3$).
@@
Appendix E User Study Details
@@
E.1 Multi-person Audio-Driven Video Generation Questionnaire
@@
E.2 Multi-concept Video Generation Questionnaire

### v2

@@ Appendix B Algorithm of Our Model Implementation
7:for $i\leftarrow 1$ to NN do
8:  $T_{i}\leftarrow\text{Rephrase}(X_{i},T)$ $\triangleright$ Generate detailed prompt via Qwen2.5-VL.
9:  $\mathbf{x}_{i}\leftarrow\text{VAE\_Encoder}(X_{i})$ $\triangleright$ Encode reference image.
10:  $\mathbf{a}_{i}\leftarrow\text{wav2vec}(Y_{i})$ $\triangleright$ Extract audio features.
@@
[…]
@@
17:   for $i\leftarrow 1$ to NN do
18:     Reference Injection: Inject $\{\mathbf{x}_{i}\}$ via concatenation and self-attention.
@@
20:     $\mathbf{Q}^{v}\leftarrow\text{Proj}_{q}(\mathbf{h}^{v})$.  $\mathbf{K}^{r}_{i},\,\mathbf{V}^{r}_{i}]\leftarrow\text{Proj}_{k,v}(\mathbf{h}^{r}_{i})$.
@@
31:     $\mathbf{Q}^{v}\leftarrow\text{Proj}_{q}(\mathbf{h}^{v})$.  $\mathbf{K}^{\text{mute}}_{i},\,\mathbf{V}^{\text{mute}}_{i}]\leftarrow\text{Proj}_{k,v}(\mathbf{a}_{i}^{\text{mute}})$.  $\mathbf{K}_{i},\,\mathbf{V}_{i}]\leftarrow\text{Proj}_{k,v}(\mathbf{a}_{i})$.
@@
Appendix C More Details of Our Model
C.1 Long Video Generation and Frame-Alignment Details
We follow sliding-window strategies (Jiang et al., 2024a; Tian et al., 2025b) by reusing several motion frames from the tail of window $w$ as the head of window w+1w\!+\!1. Reference injection is unchanged; starting frames are optional for window 1.
Frame alignment flag.
Per-frame masks have confidence scores from SAM2; frames with confidence <0.5<0.5 are excluded from the mask loss (flow-matching still applies). Videos are retained if other frames have reliable masks, so supervision remains temporally dense.
C.2 Loss Design and Audio Stack Clarifications
Focal loss.
We adopt focal loss with $\alpha=0.25,\ \gamma=2$ to mitigate class imbalance (background vs person) and stabilize early training; BCE converges more slowly and is less stable initially.
Audio features.
Audio tokens come from wav2vec 2.0; we do not add extra transformer blocks before injection. Spatial attention is unrestricted; temporally each latent attends to a local $\pm 5$ token window.
C.3 Implementation Details
@@
C.4 Ablation Details
In Tab. 4 and Fig. 5 of the paper, we validate our local audio injection designs via ablation on three variants: 1) Global audio (Lin et al., 2025a) applies audio across the entire feature map; 2) ID Embedding injects audio features with a learnable ID embedding without mask prediction; 3) Fixed Mask uses predefined static spatial masks.
For global audio, it is identical to the pretrained audio-driven model (Lin et al., 2025a). For fixed mask, it is a manually input rectangle mask upon the audio cross-attention of pretrained audio-driven model (Lin et al., 2025a). It requires the generated human being static and it needs the user to ensure there is only one person inside a mask.
For ID embedding, here we provide some implementation details. Inspired by detection transformers (Carion et al., 2020), we introduce a set of $N$ learnable ID embeddings $\mathbf{E}_{query}\in\mathbb{R}^{N\times C}$. These embeddings serve as identity tokens for the $N$ individuals. Specifically, we add the same learnable ID embedding $\mathbf{e}_{query}\in\mathbb{R}^{C}$ to a paired reference image and audio segment. In this way, we expect the model to implicitly match an individual (defined by a reference image) and audio segments in the output video to get synchronized lip movements. For the ID embedding dataset preparation, we incorporate multi-person videos into training. We use Active Speaker Detection (ASD) to extract talking persons’ identities as a bounding box with identity-number. Then we use the bounding box and SAM2 masks to match the identity-number of each timestep of audio based on their IoU. Finally we use paired audio with timestep and reference image to generate video, and use flow-matching loss only to supervised it. No mask information is provided to the model and no mask supervision is adopted.
@@ C.4 Ablation Details
Appendix D Details of Audio-driven Base Model’s Architecture and Training
This appendix provides a detailed description of the network structure and training specifics for the pretrained audio-driven single-person video generation model, OmniHuman (Lin et al., 2025a). As noted in our main paper, the text-to-video base model undergoes post-pretraining for audio conditioning, following the OmniHuman (Lin et al., 2025a) methodology. Subsequently, our multi-concept framework is built and trained on this foundation.
D.1 Audio-driven Base Model
@@
The DiT Blocks are based on the dual-stream Diffusion Transformer (DiT) (Esser et al., 2024). This transformer processes video and text tokens through multiple self-attention layers and feedforward networks (FFNs) to learn representations that are both shared and modality-specific. SwiGLU is employed as the activation function to enhance nonlinear modeling capabilities. Additionally, AdaSingle (Chen et al., 2023) is used for efficient timestep modulation. Moreover, two-thirds of the FFN weights in the deeper layers share parameters, creating a hybrid-stream design that preserves model capacity while substantially improving parameter efficiency.
D.2 Mixed Training Strategy
We utilize a multi-condition training strategy based on the framework established in OmniHuman (Lin et al., 2025a). Our method employs a two-stage progressive training scheme to develop the base model’s capabilities. In the initial stage, the model is trained solely on text-to-video (T2V) data to build fundamental video generation abilities. The second stage introduces audio-synchronized datasets to expand the model’s functionality to include audio-driven generation and reference image injection. For this second stage, we follow two core principles from OmniHuman: 1) Tasks with stronger conditioning can leverage tasks with weaker conditioning and their associated data to broaden the effective training dataset; 2) Tasks with stronger conditioning should be assigned proportionally lower training ratios. Guided by these principles, we first train the reference image injection capability before introducing the audio-driven generation objective. This staged methodology facilitates efficient knowledge transfer while ensuring stable training dynamics throughout the multi-condition learning process.
Appendix E Additional Details for Audio-driven Base Model Dataset
E.1 Dataset Curation
@@
2. Human. Utilizing the annotated video captions, we implement a rule-based system to detect keywords such as ”people”, ”human”, ”men”, ”women”, ”girl”, and ”boy”. If any of these keywords are found, the video is categorized as human-related. This technique ensures the dataset is effectively filtered to comprise only videos pertinent to human activities and interactions.
@@
4. Visual Quality. We utilize Q-align (Wu et al., 2023a) to evaluate the visual quality of the videos, filtering out clips that do not meet a predefined threshold. This procedure ensures the dataset maintains a high standard of visual clarity, which is essential for accurate analysis and robust model performance. By eliminating low-quality segments, we enhance the overall dependability and utility of the dataset.
5. Aesthetics. We employ Q-align (Wu et al., 2023a) to assess the aesthetic appeal of the videos and discard clips falling below a set threshold. Through aesthetic quality assessment, we can filter out videos containing post-production elements, thereby improving the quality of the training dataset. This measure ensures the dataset is composed of natural and unaltered video content, which better represents real-world scenarios and boosts the robustness and generalization of the trained models.
6. Motion. We use Raft (Teed & Deng, 2020) to calculate the optical flow of the videos and filter out clips exhibiting overly intense motion. This step ensures that the dataset includes only video clips with moderate and significant motion, which are more appropriate for analysis and model training. By removing clips with extreme motion, we enhance the stability and quality of the dataset, leading to more dependable results.
7. Syncnet. For audio-driven content, we employ SyncNet (Chung & Zisserman, 2017) to determine if the lip movements are synchronized with the audio. Videos displaying considerable asynchronization are excluded. This step ensures that the dataset consists only of high-quality, synchronized audio-visual data. By removing out-of-sync segments, we improve the overall quality and reliability of the dataset. Ultimately, we gather 2,000 hours of audio-driven data.
E.2 Dataset Analysis
@@
  • Human Size. Human size indicates the portion of the human body visible within the video frame. It is categorized into the following levels: Portrait (head and shoulders), Chest (from head to chest), Waist (from head to waist), Knees (from head to knees), and Full Body (the entire body is visible). To determine this, we first detect body keypoints using RTMpose (Jiang et al., 2023), and then classify the human size based on the confidence scores of these keypoints. This method ensures a robust and accurate categorization of the visible human body extent in each video.
  • Motion Amplitude. After extracting body keypoints from the video, the amplitude of human motion is computed by measuring the displacement of the chest keypoint over time, relative to the width of the shoulders. Based on these calculations, we classify motion amplitude into four categories: Slight(<0.1<0.1), Moderate(0.10.20.1-0.2), Significant(0.20.30.2-0.3), and Extreme(>0.3>0.3).
@@
Appendix F User Study Details
@@
F.1 Multi-person Audio-Driven Video Generation Questionnaire
@@
F.2 Multi-concept Video Generation Questionnaire